OpenStudy (anonymous):

Greg is 5'3 tall . The doorframe to his room is 8'2 tall . How much space is there between the top of Gregs head and the bottom of the doorframe ?

OpenStudy (anonymous):

@mathway do you mind helping me again ?

OpenStudy (mathstudent55):

You can do this 2 ways: 1. convert both measurements to inches and just subtract 2. keep the measurements in ft and in. and borrow

OpenStudy (anonymous):

I pick #2

OpenStudy (anonymous):

What should I do now @mathstudent55

Mehek (mehek14):

5ft = 60 in. 60 + 3 = 63in. 8ft = 96in. 96 + 2 = 98in. 98 - 63 = ?

OpenStudy (anonymous):

@Mehek14 it equals 35in

OpenStudy (anonymous):

Now what should I do ?

Mehek (mehek14):

2'11 2ft and 11 in.

OpenStudy (anonymous):

@Mehek14 add them together ?

OpenStudy (anonymous):

Or is 2'11my final answer ?

OpenStudy (mathstudent55):

Let's do it with borrowing.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):